OpenStudy (anonymous):

Compute Line Integral: F= <2ysinx, -xsinx>, C has parameterization r(t)= for t between 0 and 1

OpenStudy (anonymous):

\[\int\limits_{c}^{}F*dr\] in this form

OpenStudy (anonymous):

@SolomonZelman @TuringTest

OpenStudy (anonymous):

F=(2ysinx, -xsinx) C as parameterization r(t)= (t, t^2)

OpenStudy (anonymous):

@iambatman @Compassionate @abb0t

OpenStudy (anonymous):

@iGreen @eliassaab

OpenStudy (freckles):

So we need to find F(r(t)) and also r'(t) \[F(r(t))=F(x(t),y(t)) \\ \text{ where } x(t)=t \text{ and } y(t)=t^2 \\ \text{ so } F(r(t))=F(x,y)=F(t,t^2) \\ \text{ replace } x \text{ with } t \text{ and replace } y \text{ with } t^2 \\ \text{ so } F(r(t))=<2t^2\sin(t), -tsin(t)>\] Now you also need to find r'(t) from r(t)=<t,t^2>

OpenStudy (freckles):

after that we will find the dot product of F(r(t)) and r'(t)

OpenStudy (freckles):

then integrate

OpenStudy (anonymous):

alright

OpenStudy (freckles):

\[\int\limits_{0}^{1}F(r(t)) \cdot r'(t) dt \]

OpenStudy (anonymous):

so r'(t)= 1, 2t

OpenStudy (freckles):

yea

OpenStudy (anonymous):

so (1,2t)* (2t^2sint, -tsint)

OpenStudy (freckles):

\[<2t^2\sin(t),-tsin(t)> \cdot <1,2t>=?\]

OpenStudy (freckles):

yeah either way <a,b> dot <c,d>=ac+bd or <c,d> dot <a,b>=ca+db

OpenStudy (anonymous):

ok thanks

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/CalcIII/LineIntegralsVectorFields.aspx never done line integrals i was just using this as a guide :)

OpenStudy (anonymous):

so then its 2t^2sint-2t^2sint

OpenStudy (anonymous):

me neither lol

OpenStudy (freckles):

and yes exactly

OpenStudy (anonymous):

im actually taking bc the answer is 0

OpenStudy (freckles):

\[\int\limits_{0}^{1} (2t^2\sin(t)-2t^2\sin(t) ) dt \]

OpenStudy (anonymous):

oh whoops

OpenStudy (freckles):

is that right? let me look at it again

OpenStudy (anonymous):

so isnt the answer simply 0? how do you integrate 0?

OpenStudy (freckles):

integrating 0 gives you a constant because the derivative of a constant is 0 \[\int\limits_{0}^{1}0 dx=0\]

OpenStudy (freckles):

but integrating a 0 on an interval gives you 0 because you are looking at a rectangle with 0 height

OpenStudy (freckles):

i'm going to look back at this work

OpenStudy (anonymous):

yeah thats what i thought

OpenStudy (freckles):

everything we did looks right I just can't believe the answer is 0 :p

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

anyways thanks for your time

OpenStudy (anonymous):

im pretty sure the answer is correct

OpenStudy (anonymous):

anyone else got anything?

OpenStudy (freckles):

I think 0 is right. I'm just getting afraid of the answer 0 because I felt like I have already seen it so much today.

OpenStudy (turingtest):

it's been a right minute since I have done these, but I think @freckles is correct :)

OpenStudy (anonymous):

ok thanks guys, gonna close